When traveling with the current, it takes Kamal 1 hour to travel 12 miles on a boat. It takes the boat 1.5 hours to travel the same 12 miles when traveling against the current.
Found 3 solutions by Theo, Alan3354, ikleyn:
Answer by Theo(13342)


r = rate of boat (speed of boat)
t = time
d = distance.
c = rate of current (speed of current)
with the current, the formula becomes (r + c) * t = d
against the current, the formula becomes (r - c) * t = d
going with the current, the boat takes 1 hour to travel 12 miles.
therefore:
(r + c) * t = d becomes (r + c) * 1 = 12
going against the current, the boat takes 1.5 hour to travel the same 12 miles.
therefore:
(r - c) * t = d becomes (r - c) * 1.5 = 12.
your two equations that need to be solved simultancously are:
(r + c) * 1 = 12
(r - c) * 1.5 = 12
divide both sides of the second equation by 1.5 and leave the first equaion as is to get:
r + c = 12
r - c = 8
add the equations together to get:
2 * r = 20
solve for r to get:
r = 20 / 2 = 10
(r + c) * 1 = 12 becomes (10 + c) * 1 = 12
simplify to get 10 + c = 12
solve for c to get c = 12 - 10 = 2
(r - c) * 1 = 8 becomes (10 - c) * 1 = 8
simplify to get 10 - c = 8
solve for c to get c = 10 - 8 = 2
you have:
r = 10
c = 2
that's your solution.
the rate / speed of the boat is 10 miles per hour.
the rate / speed of the current is 2 miles per hour.

Answer by Alan3354(69443)

Well, that's interesting.
Do you have a question?

Answer by ikleyn(52847)

.

An educated person will  NEVER  ALLOW  HIMSELF  to post the problem to the forum (looking for help) with no question.


    - - - N E V E R - - -



It is just out of any limits of a civilized behavior.